OpenStudy (anonymous):

derivative of y=-2xsin(x^2)

OpenStudy (anonymous):

let u = x^2.

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

ok, so what are you having trouble with?

OpenStudy (anonymous):

one min ill write wat I have so far

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=-2xsin u u=x^2 derivative of (-2xsinu)= \[\frac{ dy }{ du } -2x \times \sin u + -2x \times \frac{ dy }{ du } \sin u\] derivative of x^2= 2x

OpenStudy (anonymous):

so the answer I get is (-2sin(x^2)-2xcos(x^2))(2x)

OpenStudy (anonymous):

but the answer in the book is 4x^2cos(x^2)-2sin(x^2)

OpenStudy (anonymous):

oops my bad, i thought you were integrating. anyway, i can explain it. y=-2xsin(x^2) ok so keep the constant out front -2[xsin(x^2] so take a derivative of the inside -2[x*(cos(x^2))*(2x) + sin(x^2)] -2[2x^2*cos(x^2) + sin(x^2)] -4x^2cos(x^2) -2sin(x^2)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you're welcome